# Distributed Multiple Fault Detection and Estimation in DC Microgrids with Unknown Power Loads

### v1

@@ 3. Estimation of Actuator Faults
In this section, we provide the design method for $\mathcal{D}_{a,i}$. A fault estimation filter is developed as the core solution. To facilitate the filter design, the state-space model (4) is reformulated into the DAE form:
@@
where the polynomial row vector $N_{i}(p)=\sum^{d_{N}}_{j=0}p^{j}N_{i,j}$, $N_{i,j}\in{\mathbb{R}}^{1\times 6}$ and $d_{N}$ is the degree of $N_{i}(p)$. The denominator $a(p)$ is a polynomial defined as $a(p)=p^{d_{a}}+\sum^{d_{a}-1}_{j=0}p^{j}a_{j}$, where $d_{a}$ denotes the degree of $a(p)$ and the coefficients $a_{j}\in{\mathbb{R}}$. For simplicity of filter design, $a(p)$ is given with all roots lying in the left-half plane and is identical for all $\mathcal{D}_{a,i}$. Its degree is set as $d_{a}=d_{N}+1$ to ensure that the filter is strictly proper.
@@
Recall the first design requirement for $\mathcal{D}_{a,i}$, which stipulates that $P_{i}$ and $f_{I,i}$, have no influence on $\hat{f}_{a,i}$. To achieve this, we introduce the following condition
@@
Based on (7), the design requirements for $\mathcal{D}_{a,i}$ have been translated into constraints (8a)-(8c) on the mapping relations. In the following proposition, we further formulate a tractable optimization problem for solving the parameters of $N_{i}(p)$ based on (8a)-(8c). Before proceeding, the observable canonical form of $N_{i}(p)/a(p)$, denoted by $\{A_{r},B_{r,i},C_{r}\}$, is provided to facilitate computation of the $\mathcal{H}_{2}$ norm. Here, $A_{r}$ is given by the coefficients of $a(p)$, which we keep the same for all $i$, while $B_{r,i}$ is different for different $i$. Specifically, the matrices are given by
@@
The design approach of the actuator fault estimation filter is then presented in the following proposition.
@@
The proof is similar to that of Theorem 3.1 from [33] and thus is omitted here. ∎
The feasibility of (10) depends on the rank properties of $\bar{H}_{i}$ and $\mathcal{E}_{i}$. Specifically, the condition for (10a) to have non-trivial solutions is that $\bar{H_{i}}$ does not have full row rank. This can be satisfied by selecting the numerator degree of the filter $d_{N}$ such that $6(d_{N}+1)>\text{rank}(\bar{H}_{i})$, where $6(d_{N}+1)$ is the row number of $\bar{H_{i}}$. Meanwhile, since all roots of $a(p)$ are in the left-half plane, $a(0)\neq 0$ is guaranteed. Consequently, (10b) requires that $\mathcal{E}_{i}$ lies outside the column range space of $\bar{H}_{i}$, i.e., $\text{rank}([\bar{H_{i}}~\mathcal{E}_{i}])>\text{rank}(\bar{H_{i}})$. Otherwise, a feasible solution to $N_{i,0}\bar{H}_{i}=0$ leads to $N_{i,0}\mathcal{E}_{i}=0$ as well, and (10b) is not satisfied.
@@ 4. Estimation of aggregate faulty line current
In this section, we present the design method for the diagnosis block $\mathcal{D}_{l,i}$, which estimates the aggregate faulty line current $f_{I,i}$ induced by faults on power lines connected to DG unit $i$. As shown in Fig. 3, $\mathcal{D}_{l,i}$ consists of four sub-blocks:
@@
    A set of line current estimators to estimate the fault-free part of the total line currents $\sum_{k=1}^{m}\mathbb{B}_{ik}I_{k}$ injected into DG unit $i$, which produces $\sum_{k=1}^{m}\mathbb{B}_{ik}\hat{I}_{k,h}$.
@@ 4.1. Design of Pre-filter 11
where $\mathcal{N}_{i}(p)=\sum^{d_{\mathcal{N}}}_{j=0}p^{j}\mathcal{N}_{i,j}$ with design parameters $\mathcal{N}_{i,j}\in{\mathbb{R}}^{1\times 6}$ and degree $d_{\mathcal{N}}$. The denominator $a(p)$ is the same as that in (6). From (11) and (12), $r_{i}$ also equals to
@@ 4.3. Design of Pre-filter 22
where the input consists of the aggregate estimates of fault-free line currents, the output $\hat{r}_{I,i}$ represents the estimated contribution of $\sum_{k=1}^{m}\mathbb{B}_{ik}I_{k,h}$ to $r_{i}$. According to (13), (15), and (16), the residual $\tilde{r}_{i}$ in Fig. 3 encapsulating the information of $f_{I,i}$ and $P_{i}$ is then generated by subtracting $\hat{r}_{I,i}$ from $r_{i}$, yielding
@@
$\left\{\begin{array}[]{l}\dot{x}_{\tilde{r}_{i}}(t)=A_{r}x_{\tilde{r}_{i}}(t)+B_{\mathcal{G},i}\left(f_{I,i}(t)+\frac{P_{i}(t)}{V_{i}(t)}\right)+B_{\varpi,i}\varpi_{i}(t),\\ \tilde{r}_{i}(t)=C_{r}x_{\tilde{r}_{i}}(t),\end{array}\right$. (20)
@@
$B_{\mathcal{G},i}=\begin{bmatrix}\mathcal{N}^{\top}_{i,0}&\mathcal{N}^{\top}_{i,1}&\dots&\mathcal{N}^{\top}_{i,d_{\mathcal{N}}}\end{bmatrix}^{\top}\mathcal{G}_{i},~B_{\varpi,i}=\begin{bmatrix}\mathcal{N}^{\top}_{i,0}&\mathcal{N}^{\top}_{i,1}&\dots&\mathcal{N}^{\top}_{i,d_{\mathcal{N}}}\end{bmatrix}^{\top}\begin{bmatrix}\mathcal{G}_{i}&{\bf I}_{6}\end{bmatrix}$.
@@ 4.4. Design of Faulty Line Current Estimator
The three sub-blocks designed above serve as preparatory steps for estimating $f_{I,i}$. As illustrated in Fig. 3, the derived $\tilde{r}_{i}$ is processed by the estimator in the final sub-block to compute $\hat{f}_{I,i}$. In this final stage, a discrete-time estimation approach is developed, and thus the state-space model (20) is discretized with a sampling period $t_{s}$. For consistency and simplicity, the system matrix notations of the discrete-time model remain unchanged. To address the potential ill-posedness issue in estimating $f_{I,i}$ mentioned in the problem statement session, the proposed approach proceeds through two sequential phases:
@@
    Detection and differentiation phase. Before any power line faults, i.e., $f_{I,i}=0$, a residual $\tilde{\Upsilon}_{i,T}$ derived from the estimate of $P_{i}$ is introduced to detect both step load changes and power line faults. The discrimination between the two events is further achieved by leveraging their unique transient signatures in the residual signal $\tilde{\Upsilon}_{i,T}$.
@@
    Estimation phase. Once a power line fault is detected, a dedicated estimator is activated to estimate $f_{I,i}$.
@@ 4.4.1. Detection and differentiation of power line faults and step load changes
The proposed detection and differentiation approach relies on power load estimation before faults occur on the power lines connected to DG unit $i$. Suppose that $f_{I,i}=0$ and $P_{i}$ is constant initially. To obtain power load estimation, we construct the following parity-space relation over a sliding window $T$ based on (20):
@@
To simplify the subsequent analysis, we further introduce the orthogonal projection of $\mathcal{O}_{i,T}$, i.e., $\mathcal{O}^{\bot}_{i,T}=I-\mathcal{O}_{i,T}\mathcal{O}^{{\dagger}}_{i,T}$. Multiplying both sides of (21) from the left by $\mathcal{O}^{\bot}_{i,T}$ leads to
@@
where $\Upsilon_{i,T}(k)=\mathcal{O}^{\bot}_{i,T}\tilde{\bm{r}}_{i,T}(k)\in{\mathbb{R}}^{n_{\Upsilon}}$, $\Psi_{i,T}(k-1)=\mathcal{O}^{\bot}_{i,T}\mathcal{Z}_{i_{1},T}\bm{\mathcal{V}}_{i,T-1}(k-1)$, and $\Omega_{i,T}(k-1)=\mathcal{O}^{\bot}_{i,T}\mathcal{Z}_{i_{2},T}\bm{\mathcal{\varpi}}_{i,T-1}(k-1)$. For notational simplicity, the time index is omitted hereafter, e.g., $\Upsilon_{i,T}(k)\rightarrow\Upsilon_{i,T}$.
Given (22), the weighted least-squares method is employed here to estimate $P_{i}$. The estimated value $\hat{P}_{i}$ is given by
@@
where $\Phi_{i,T}=(\Psi^{\top}_{i,T}\Sigma^{-1}_{\Omega_{i,T}}\Psi_{i,T})^{-1}\Psi^{\top}_{i,T}\Sigma^{-1}_{\Omega_{i,T}}$ and $\Sigma_{\Omega_{i,T}}=\mathcal{O}^{\bot}_{i,T}\mathcal{Z}_{i_{2},T}\Sigma_{\bm{\mathcal{\varpi}}_{i,T-1}}(\mathcal{O}^{\bot}_{i,T}\mathcal{Z}_{i_{2},T})^{\top}\succ 0$ represents the covariance matrix of $\Omega_{i,T}$.
To detect load changes or power line faults, the following residual is introduced:
$\tilde{\Upsilon}_{i,T}=\Upsilon_{i,T}-\Psi_{i,T}\hat{P}_{l_{i}},$ (24)
where the $\kappa$-th row of $\tilde{\Upsilon}_{i,T}$ is $\tilde{\Upsilon}^{[\kappa]}_{i,T}=\Upsilon^{[\kappa]}_{i,T}-\Psi^{[\kappa]}_{i,T}\hat{P}_{i},~\kappa\in\{1,\dots,n_{\Upsilon}\}$. Note that when there are no load changes and line faults, $\hat{P}_{i}$ obtained from (23) is the minimum-variance unbiased estimate of $P_{i}$ [34, Section 4.5]. Therefore, we have
@@
where $e_{\kappa}$ is a row vector with all entries zero except for the $\kappa$-th entry being $1$. According to Chebyshev’s inequality, in the absence of load changes and power line faults, the probability that $\tilde{\Upsilon}^{[\kappa]}_{i,T}(k)$ lies inside the threshold interval $\varepsilon^{[\kappa]}_{i}(k),\varepsilon^{[\kappa]}_{i}(k)]$ satisfies:
@@
    Step load changes in $P_{i}$. Entries in $\tilde{\Upsilon}_{i,T}$ can temporarily exceed the threshold interval $\varepsilon^{[\kappa]}_{i}(k),\varepsilon^{[\kappa]}_{i}(k)]$ under the effect of step load changes. Then, with probability greater than $1-1/\alpha^{2}$, the entries return to and remain within the threshold bounds after $T$ steps following the load changes.
@@
    and the faults can be distinguished from step changes in $P_{i}$ if there consistently exists at least one entry in $\tilde{\Upsilon}_{i,T}(k)$ that satisfies $\tilde{\Upsilon}^{[\kappa]}_{i,T}(k)\notin[-\varepsilon^{[\kappa]}_{i}(k),\varepsilon^{[\kappa]}_{i}(k)]$ for at least $T$ consecutive steps.
@@
According to Proposition 4.1, step load changes and power line faults can be distinguished by evaluating the duration of their effects on $\tilde{\Upsilon}_{i,T}$. To calculate the duration time during which not all $\tilde{\Upsilon}^{[\kappa]}_{i,T}$ remain within the threshold intervals, we introduce $\mathcal{T}_{i}(k)$ to record the most recent time instant at which all $\tilde{\Upsilon}^{[\kappa]}_{i,T}$ are inside the thresholds, which is
@@
    Step load changes or line faults ($\sigma(k)=1$). If $T/2\leq k-\mathcal{T}_{i}(k)<T$, then there exists at least one element $\tilde{\Upsilon}^{[\kappa]}_{i,T}(t)\notin[-\varepsilon^{[\kappa]}_{i}(t),\varepsilon^{[\kappa]}_{i}(t)]$ for each $t\in\{\mathcal{T}_{i}(k)+1,\dots,k\}$ with duration less than $T$ while larger than and equal to $T/2$;
@@
    Persistent line faults ($\sigma(k)=2$). If $k-\mathcal{T}_{i}(k)\geq T$, then there exists $\kappa\in\{1,\dots,n_{\Upsilon}\}$ such that $\tilde{\Upsilon}^{[\kappa]}_{i,T}(k)\notin[-\varepsilon^{[\kappa]}_{i}(k),\varepsilon^{[\kappa]}_{i}(k)]$ for at least $T$ consecutive steps.

### v2

@@ 3. Estimation of Actuator Faults
In this section, we provide the design method for $\mathcal{D}_{a,i}$, where a fault estimation filter is developed as the core solution. To facilitate the filter design, the state-space model (4) is reformulated into the DAE form:
@@
where the polynomial row vector $N_{i}(p)=\sum^{d_{N}}_{j=0}p^{j}N_{i,j}$, $N_{i,j}\in{\mathbb{R}}^{1\times 6}$ and $d_{N}$ is the degree of $N_{i}(p)$. The denominator $a(p)$ is a polynomial defined as $a(p)=p^{d_{a}}+\sum^{d_{a}-1}_{j=0}p^{j}a_{j}$, where the coefficient $a_{j}\in{\mathbb{R}}$ and $d_{a}$ denotes the degree of $a(p)$. For simplicity of filter design, $a(p)$ is given with all roots lying in the left-half plane and is identical for all $\mathcal{D}_{a,i}$. Its degree is set as $d_{a}=d_{N}+1$ to ensure that the filter is strictly proper.
@@
Recall the first design requirement for $\mathcal{D}_{a,i}$, which stipulates that $P_{i}$ and $f_{I,i}$ have no influence on $\hat{f}_{a,i}$. To achieve this, we introduce the following condition
@@
The design requirements for $\mathcal{D}_{a,i}$ have been translated into constraints (8a)-(8c) on the mapping relations. In the following proposition, we further formulate a tractable optimization problem for solving the parameters of $N_{i}(p)$ based on (8a)-(8c). Before proceeding, the observable canonical form of $N_{i}(p)/a(p)$, denoted by $\{A_{r},B_{r,i},C_{r}\}$, is provided to facilitate computation of the $\mathcal{H}_{2}$ norm. Here, $A_{r}$ is given by the coefficients of $a(p)$, which we keep the same for all $i$, while $B_{r,i}$ is different for different $i$. Specifically, the matrices are given by
@@
The design approach of the actuator fault estimation filter (6) is then presented in the following proposition.
@@
The proof is similar to that of Theorem 3.1 from [6], and only a brief sketch is provided. First, based on the polynomial multiplication rule, condition (10a) enforces the coefficients of the product $N_{i}(p)H_{i}(p)$ to be zero, which guarantees that condition (8a) is satisfied. Second, condition (10b) is obtained by setting the operator $p=0$ in condition (8c), yielding the steady-state gain of the corresponding transfer function. Finally, the linear matrix inequalities in (10c) follow from classical results on the $\mathcal{H}_{2}$ norm, which ensures the satisfaction of (8b). This completes the proof. ∎
It is worth noting that there always exist feasible filters satisfying (8a)-(8c). First, conditions (8a) and (8c) can be satisfied simultaneously because, from the structures of $D_{i}$ and $E_{i}$, the fault direction does not lie in the disturbance subspace, thereby guaranteeing disturbance decoupling and fault sensitivity [5, Theorem 6.2]. Second, condition (8b) requires the corresponding transfer function to be stable and strictly proper so that its $\mathcal{H}_{2}$ norm is finite, which can be ensured by an appropriate choice of $a(p)$. In addition, the feasibility of problem (10) depends on the rank conditions of $\bar{H}_{i}$ and $\mathcal{E}_{i}$. Specifically, (10a) admits nontrivial solutions only if $\bar{H_{i}}$ is not of full row rank, which can be ensured by selecting $d_{N}$ such that $6(d_{N}+1)>\text{rank}(\bar{H}_{i})$. Meanwhile, since all roots of $a(p)$ are in the left-half plane, $a(0)\neq 0$ is guaranteed. Consequently, condition (10b) requires that $\mathcal{E}_{i}$ lies outside the column range space of $\bar{H}_{i}$, i.e., $\text{rank}([\bar{H_{i}}~\mathcal{E}_{i}])>\text{rank}(\bar{H_{i}})$. Otherwise, any solution satisfying $N_{i,0}\bar{H}_{i}=0$ leads to $N_{i,0}\mathcal{E}_{i}=0$ as well, and (10b) can not be satisfied.
Remark 3.2 (Differences with previous work).
In our previous work [7], disturbances were non-decouplable, and the filter design was formulated as a quadratic programming problem with a fixed denominator. In contrast, this work (i) exploits structural disturbance decouplability via (10a), (ii) incorporates stochastic noise via the $\mathcal{H}_{2}$-norm method, and (iii) allows joint numerator–denominator design since the parameters of $a(p)$ are embedded in $A_{r}$, leading to a bi-linear formulation when the denominator is also treated as a design variable.
@@ 4. Estimation of aggregate faulty line current
In this section, we present the design method for the diagnosis block $\mathcal{D}_{l,i}$, which estimates the aggregate faulty line current $f_{I,i}$ induced by faults on power lines connected to DG unit $i$ under the piecewise-constant power load assumption. As shown in Fig. 3, $\mathcal{D}_{l,i}$ consists of four sub-blocks:
@@
    A set of line current estimators to estimate the fault-free part of the total line currents $\sum_{k=1}^{m}\mathbb{B}_{ik}I_{k}$ injected into DG unit $i$, and produces $\sum_{k=1}^{m}\mathbb{B}_{ik}\hat{I}_{k,h}$.
@@ 4.1. Design of Pre-filter 11
where $\mathcal{N}_{i}(p)=\sum^{d_{\mathcal{N}}}_{j=0}p^{j}\mathcal{N}_{i,j}$ with design parameters $\mathcal{N}_{i,j}\in{\mathbb{R}}^{1\times 6}$ and degree $d_{\mathcal{N}}$. The denominator $a(p)$ is the same as that in (6). From (11) and (12), $r_{i}$ also equals
@@ 4.3. Design of Pre-filter 22
where the input is an estimate of the aggregate fault-free line currents, the output $\hat{r}_{I,i}$ represents the estimated contribution of $\sum_{k=1}^{m}\mathbb{B}_{ik}I_{k,h}$ to $r_{i}$. According to (13), (15), and (16), the residual $\tilde{r}_{i}$ in Fig. 3 encapsulating the information of $f_{I,i}$ and $P_{i}$ is then generated by subtracting $\hat{r}_{I,i}$ from $r_{i}$, yielding
@@
$\dot{x}_{\tilde{r}_{i}}(t)=A_{r}x_{\tilde{r}_{i}}(t)+B_{\mathcal{G},i}\left(f_{I,i}(t)+\frac{P_{i}(t)}{V_{i}(t)}\right)+B_{\varpi,i}\varpi_{i}(t),$
$\tilde{r}_{i}(t)=C_{r}x_{\tilde{r}_{i}}(t),$ (18)
@@
$B_{\mathcal{G},i}=\begin{bmatrix}\mathcal{N}^{\top}_{i,0}&\mathcal{N}^{\top}_{i,1}&\dots&\mathcal{N}^{\top}_{i,d_{\mathcal{N}}}\end{bmatrix}^{\top}\mathcal{G}_{i},$
$B_{\varpi,i}=\begin{bmatrix}\mathcal{N}^{\top}_{i,0}&\mathcal{N}^{\top}_{i,1}&\dots&\mathcal{N}^{\top}_{i,d_{\mathcal{N}}}\end{bmatrix}^{\top}\begin{bmatrix}\mathcal{G}_{i}&{\bf I}_{6}\end{bmatrix}$.
@@ 4.4. Design of Faulty Line Current Estimator
The three sub-blocks designed above serve as preparatory steps for estimating $f_{I,i}$. As illustrated in Fig. 3, the derived $\tilde{r}_{i}$ is processed by the estimator in the final sub-block to compute $\hat{f}_{I,i}$. In this final stage, a discrete-time estimation approach is developed, and thus the state-space model (4.3) is discretized with a sampling period $t_{s}$. For consistency and simplicity, the system matrix notations of the discrete-time model remain unchanged. To address the potential ill-posedness issue in estimating $f_{I,i}$ mentioned in the problem statement session, the proposed approach proceeds through two sequential phases:
@@
    Detection and differentiation phase. Before the occurrence of any power line faults, i.e., $f_{I,i}=0$, a residual $\tilde{\Upsilon}_{i,T}$ derived from the estimate of $P_{i}$ is introduced to detect both step load changes and power line faults. The discrimination between the two events is further achieved by leveraging their distinct transient behaviors in $\tilde{\Upsilon}_{i,T}$, in particular, the duration of threshold violation.
@@
    Estimation phase. Once a power line fault is detected, a dedicated regularized LS estimator is activated to estimate $f_{I,i}$.
@@ 4.4.1. Detection and differentiation of power line faults and step load changes
The proposed detection and differentiation approach relies on power load estimation before faults occur on the power lines connected to DG unit $i$. To this end, suppose that $f_{I,i}=0$ and $P_{i}$ is constant initially based on Assumption 2.4. To obtain power load estimation, we construct the following parity-space relation over a sliding window $T$ based on (4.3):
@@
To eliminate the effect of unknown state $x_{\tilde{r}_{i}}(k-T+1)$ on the residual, we further introduce the orthogonal projection of $\mathcal{O}_{i,T}$, i.e., $\mathcal{O}^{\bot}_{i,T}=I-\mathcal{O}_{i,T}\mathcal{O}^{{\dagger}}_{i,T}$. Multiplying both sides of (19) from the left by $\mathcal{O}^{\bot}_{i,T}$ leads to
@@
where $\Upsilon_{i,T}(k)=\mathcal{O}^{\bot}_{i,T}\tilde{\bm{r}}_{i,T}(k)\in{\mathbb{R}}^{n_{\Upsilon}}$,
$\Psi_{i,T}(k-1)=\mathcal{O}^{\bot}_{i,T}\mathcal{Z}_{i_{1},T}\bm{\mathcal{V}}_{i,T-1}(k-1),$
$\Omega_{i,T}(k-1)=\mathcal{O}^{\bot}_{i,T}\mathcal{Z}_{i_{2},T}\bm{\mathcal{\varpi}}_{i,T-1}(k-1).$
For notational simplicity, the time index is omitted hereafter, e.g., $\Upsilon_{i,T}(k)\rightarrow\Upsilon_{i,T}$.
Given (20), the weighted LS method is employed here to estimate $P_{i}$. The estimated value $\hat{P}_{i}$ is given by
@@
where $\Phi_{i,T}=(\Psi^{\top}_{i,T}\Sigma^{-1}_{\Omega_{i,T}}\Psi_{i,T})^{-1}\Psi^{\top}_{i,T}\Sigma^{-1}_{\Omega_{i,T}}$ and $\Sigma_{\Omega_{i,T}}=\mathcal{O}^{\bot}_{i,T}\mathcal{Z}_{i_{2},T}\Sigma_{\bm{\mathcal{\varpi}}_{i,T-1}}(\mathcal{O}^{\bot}_{i,T}\mathcal{Z}_{i_{2},T})^{\top}\succ 0$ represents the covariance matrix of $\Omega_{i,T}$. To detect load changes and power line faults, the following residual is introduced:
$\tilde{\Upsilon}_{i,T}=\Upsilon_{i,T}-\Psi_{i,T}\hat{P}_{i},$ (22)
which characterizes the mismatch between $\Upsilon_{i,T}$ and its reconstruction based on $\hat{P}_{i}$. In the absence of load changes and line faults, $\hat{P}_{i}$ obtained from (21) is the minimum-variance unbiased estimate of $P_{i}$ [14, Section 4.5], and  $\tilde{\Upsilon}_{i,T}$ fluctuates around 0 due to noise. When a load change or power line fault occurs, the mismatch between the actual and estimated loads causes $\tilde{\Upsilon}_{i,T}$ to deviate from 0 apparently, thereby indicating the occurrence of load variations or faults.
We further define the $\kappa$-th row of $\tilde{\Upsilon}_{i,T}$ as $\tilde{\Upsilon}^{[\kappa]}_{i,T}=\Upsilon^{[\kappa]}_{i,T}-\Psi^{[\kappa]}_{i,T}\hat{P}_{i},~\kappa\in\{1,\dots,n_{\Upsilon}\}$. When there are no load changes and line faults, according to [14, Section 4.5], it holds that
@@
where $e_{\kappa}$ is a row vector with all entries zero except for the $\kappa$-th entry being $1$. Based on Chebyshev’s inequality, the probability that $\tilde{\Upsilon}^{[\kappa]}_{i,T}(k)$ lies inside the threshold interval $\varepsilon^{[\kappa]}_{i}(k),\varepsilon^{[\kappa]}_{i}(k)]$ satisfies:
@@
    Step load changes in $P_{i}$. Entries in $\tilde{\Upsilon}_{i,T}$ can temporarily exceed the threshold interval $\varepsilon^{[\kappa]}_{i}(k),\varepsilon^{[\kappa]}_{i}(k)]$ due to a step load change. Then, with probability greater than $1-1/\alpha^{2}$, the entries return to and remain within the threshold bounds within $T$ steps after the load changes.
@@
    which remains nonzero as long as the fault persists. Then, power line faults can be distinguished from step changes in $P_{i}$ if there consistently exists at least one entry in $\tilde{\Upsilon}_{i,T}(k)$ that satisfies $\tilde{\Upsilon}^{[\kappa]}_{i,T}(k)\notin[-\varepsilon^{[\kappa]}_{i}(k),\varepsilon^{[\kappa]}_{i}(k)]$ for at least $T$ consecutive steps.
Therefore, by monitoring whether threshold violations are transient (disappearing within $T$ steps) or persistent (lasting at least $T$ consecutive steps), step load changes and line faults can be discriminated over a time window.
@@
To calculate the duration time during which not all $\tilde{\Upsilon}^{[\kappa]}_{i,T}$ remain within the threshold intervals, we introduce $\mathcal{T}_{i}(k)$ to record the most recent time instant at which all $\tilde{\Upsilon}^{[\kappa]}_{i,T}$ are inside the thresholds, which is
@@
    Step load changes or line faults ($\sigma(k)=1$). If $T/2\leq k-\mathcal{T}_{i}(k)<T$, then there exists at least one element $\tilde{\Upsilon}^{[\kappa]}_{i,T}(k^{\prime})\notin[-\varepsilon^{[\kappa]}_{i}(k^{\prime}),\varepsilon^{[\kappa]}_{i}(k^{\prime})]$ for each $k^{\prime}\in\{\mathcal{T}_{i}(k)+1,\dots,k\}$ with duration less than $T$ but greater than or equal to $T/2$. This corresponds to an intermediate state, in which an event has been detected but not distinguished yet. The discrimination result is obtained by further monitoring whether the threshold violation persists up to $T$ steps or not;
@@
    Persistent line faults ($\sigma(k)=2$). If $k-\mathcal{T}_{i}(k)\geq T$, then there exists $\kappa\in\{1,\dots,n_{\Upsilon}\}$ such that $\tilde{\Upsilon}^{[\kappa]}_{i,T}(k)\notin[-\varepsilon^{[\kappa]}_{i}(k),\varepsilon^{[\kappa]}_{i}(k)]$ for at least $T$ consecutive steps, indicating the occurrence of a line fault.
@@
Note that the differentiation strategy is developed under the piecewise-constant power load assumption. However, load generally exhibits non-ideal step dynamics in practice, such as start-up transients of electric-vehicle chargers. Nonetheless, the proposed diagnosis approach remains applicable, although the theoretical guarantees established in Proposition 4.1, such as the transient effect on the residual vanishing within $T$ steps, no longer strictly hold. In particular, slowly varying load transients may not trigger the detection threshold as $\Delta P$ is small, while fast transients are more likely to be detected since they resemble step-like changes. In both cases, the proposed line-fault estimation performance is maintained, which has been demonstrated by the simulation provided in the Appendix.
Remark 4.2 (Selection of $T$).
There are some principles in selecting $T$. First, $T$ must satisfy $T>d_{\mathcal{N}}+1$ to ensure the existence of the orthogonal complement of $\mathcal{O}_{i,T}$ in (20). Second, $T$ should not be chosen too short, since slowly varying faulty line currents may then appear approximately constant over the window and become difficult to detect because the effect is not sufficiently accumulated over that time interval. On the other hand, a larger $T$ generally improves the detectability of slowly varying faults and robustness to noise, but also increases diagnosis delay and computational burden. Therefore, $T$ should be chosen as a compromise among detectability, delay, and complexity.